# Compressing local vertex functions from the multipoint numerical renormalization group using quantics tensor cross interpolation

### v2

@@ III.1 Single impurity Anderson model
The resulting rank $\chi$ is compared with the ranks $\chi_{\boldsymbol{k}}$ of individual Keldysh components in Tab. 2. We observe the ratio $\chi^{4}/\max_{\boldsymbol{k}}\chi_{\boldsymbol{k}}^{4}\approx 21.77$, thus MPO-MPO contractions take about 22 times longer for two entire vertices than for two individual components. On the other hand, the latter type of contraction would have to be performed $255^{2}=$ 25 times. However, Tab. 2 shows that some Keldysh components have a significantly lower bond dimension than $\max_{\boldsymbol{k}}\chi_{\boldsymbol{k}}$. In summary, both approaches are worth investigating, and a conclusive comparison is only possible in the context of a specific QTT-based diagrammatic code.
@@ Acknowledgements.
We thank Seung-Sup Lee and Jae-Mo Lihm for helpful discussions. We thank Seung-Sup Lee for providing the mpNRG code used for computing the PSFs. This work was funded in part by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). It is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. We further acknowledge support from the DFG grant LE 3883/2-2. We gratefully acknowledge computational resources from grant INST 86/1885-1 FUGG of the German Research Foundation (DFG). MF, AG, MR, and NR acknowledge support from the International Max-Planck Research School for Quantum Science and Technology (IMPRS-QST). NR acknowledges funding from the Studienstiftung des deutschen Volkes and the Marianne-Plehn-Programm of the state of Bavaria.
@@ Appendix A Computing broadened Keldysh kernels
The broadening parameters $\gamma_{\mathrm{L}}$ and $\sigma_{\mathrm{sLG}}$ along with other numerical settings for all physical parameter sets are listed in Tab. 3. However, the linear broadening $\gamma_{L}$ is multiplied with a prefactor that depends on the current permutation $p$, the fully retarded index $\lambda$ and the dimension $i$. This scheme will be explained further at the end of this section (see Eq. (38)). The integral (35a) is performed by trapezoidal quadrature, where $\omega^{\prime}$ and $\epsilon^{\prime}$ are discretized on logarithmic grids. These grids contain 00, are symmetric around the origin and range from $e_{\min}$ to $e_{\max}$ with $e_{\mathrm{step}}$ points per decade (see Tab. 3). For the $\epsilon^{\prime}$ grid, $e_{\min}$ is automatically replaced by a lower boundary $x_{\min}$ with $0<x_{\min}<e_{\min}$ if the low-frequency tail of the log-Gaussian broadening kernel extends below $e_{\min}$. This ensures an accurate integration of $\delta_{\mathrm{sLG}}(\epsilon^{\prime},\epsilon)$. The energies $\epsilon$ specifying the location of the spectral function peaks also reside on a logarithmic grid, which arises from the mpNRG computation.
@@
where $\mathcal{P}$ denotes the Cauchy principal value (PV) integral. Also, recall the definition (10c) of $\gamma^{\lambda}_{0,i}$. Importantly, $\delta_{b}(\omega^{\prime},\epsilon)$ has been computed on a logarithmic grid, while $\omega$ in the broadened Keldysh kernel $K_{b}^{[\lambda]}(\omega,\epsilon)$ defined in Eq. (10b) resides on a linear grid. This is because the external frequency grids on which we compute vertices are also linear. To obtain the imaginary part of Eq. (LABEL:eq:app_KramersKronig) on the linear grid, we use linear interpolation of $\delta_{b}(\omega,\epsilon)$ in the argument $\omega$. Computing the real part, i.e., the PV integral is slightly more involved: By the linear interpolation performed for the imaginary part, $\delta_{b}(\omega,\epsilon)$ can be viewed as a piecewise linear function. We split the PV integral over $\delta_{b}(\omega^{\prime},\epsilon)$ into PV integrals over linear functions $(a_{i,\epsilon}\omega^{\prime}+b_{i,\epsilon})$ on intervals $[\omega_{i}^{\prime},\omega_{i+1}^{\prime}]$. These are evaluated using the formula
@@
The sum over all PV integrals of the form (37) then yields the real part of the broadened Keldysh kernel $K_{b}(\omega,\epsilon)$. For this scheme to be accurate, the extent $[-e_{\max},e_{\max}]$ of the logarithmic $\omega^{\prime}$ grid should be significantly larger than the frequency box delimited by $\omega_{\max}$. Comparing the values of $e_{\max}$ given in Tab. 3 with our default frequency box size $\omega_{\max}=0.65$, one verifies that this is the case.
@@
$\gamma_{\mathrm{L},i}^{\lambda}=\begin{cases}\gamma_{\mathrm{L}}\cdot(\ell-i)\quad\mathrm{for}\quad i\geq\lambda,\\ \gamma_{\mathrm{L}}\cdot i\quad\mathrm{for}\quad i<\lambda.\\ \end{cases}$ (38)
@@ Appendix B Compression of 1D and 2D vertex contributions
$P^{k_{1}k_{2}(k_{3}+k_{4})}=\begin{cases}\frac{1}{\sqrt{2}}\quad&\text{if }\sum_{i}k_{i}\text{ is odd},\\ \hfil 0\quad&\text{else},\\ \end{cases}$ (40)
@@ Appendix C Compression for different channels, flavors, Keldysh components
[…]

### v3

@@ III.1 Single impurity Anderson model
The resulting rank $\chi$ is compared with the ranks $\chi_{\boldsymbol{k}}$ of individual Keldysh components in Tab. 2. We observe the ratio $\chi^{4}/\max_{\boldsymbol{k}}\chi_{\boldsymbol{k}}^{4}\approx 21.77$, thus MPO-MPO contractions take about 22 times longer for two entire vertices than for two individual components. On the other hand, the latter type of contraction would have to be performed $5^{2}=25$ times. However, Tab. 2 shows that some Keldysh components have a significantly lower bond dimension than $\max_{\boldsymbol{k}}\chi_{\boldsymbol{k}}$. In summary, both approaches are worth investigating, and a conclusive comparison is only possible in the context of a specific QTT-based diagrammatic code.
@@ Acknowledgements.
We thank Seung-Sup Lee and Jae-Mo Lihm for helpful discussions. We thank Seung-Sup Lee for providing the mpNRG code used for computing the PSFs. This work was funded in part by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy EXC-2111 (Project No. 390814868). It is part of the Munich Quantum Valley, supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. We further acknowledge support from the DFG grant LE 3883/2-2. We gratefully acknowledge computational resources from grant INST 86/1885-1 FUGG of the German Research Foundation (DFG) and from the GCS Supercomputer SuperMUC-NG at the Leibniz Supercomputing Centre in Munich provided by the Gauss Centre for Supercomputing e.V. MF, AG, MR, and NR acknowledge support from the International Max-Planck Research School for Quantum Science and Technology (IMPRS-QST). NR acknowledges funding from the Studienstiftung des deutschen Volkes and the Marianne-Plehn-Programm of the state of Bavaria.
@@ Appendix A Computing broadened Keldysh kernels
The broadening parameters $\gamma_{\mathrm{L}}$ and $\sigma_{\mathrm{sLG}}$ along with other numerical settings for all physical parameter sets are listed in Tab. 3. However, the linear broadening $\gamma_{L}$ is multiplied with a prefactor that depends on the current permutation $p$, the fully retarded index $\lambda$ and the dimension $i$. This scheme will be explained further at the end of this section (see Eq. (38)). The integral (35a) is performed by trapezoidal quadrature, where $\omega^{\prime}$ and $\epsilon^{\prime}$ are discretized on logarithmic grids. These grids contain 0, are symmetric around the origin and range from $e_{\min}$ to $e_{\max}$ with $e_{\mathrm{step}}$ points per decade (see Tab. 3). For the $\epsilon^{\prime}$ grid, $e_{\min}$ is automatically replaced by a lower boundary $x_{\min}$ with $0<x_{\min}<e_{\min}$ if the low-frequency tail of the log-Gaussian broadening kernel extends below $e_{\min}$. This ensures an accurate integration of $\delta_{\mathrm{sLG}}(\epsilon^{\prime},\epsilon)$. The energies $\epsilon$ specifying the location of the spectral function peaks also reside on a logarithmic grid, which arises from the mpNRG computation.
@@
where $\mathcal{P}$ denotes the Cauchy principal value (PV) integral. Also, recall the definition (10c) of $\gamma^{\lambda}_{0,i}$. Importantly, $\delta_{b}(\omega^{\prime},\epsilon)$ has been computed on a logarithmic grid, while $\omega$ in the broadened Keldysh kernel $K_{b}^{[\lambda]}(\omega,\epsilon)$ defined in Eq. (10b) resides on a linear grid. This is because the external frequency grids on which we compute vertices are also linear. To obtain the imaginary part of Eq. (LABEL:eq:app_KramersKronig) on the linear grid, we use linear interpolation of $\delta_{b}(\omega,\epsilon)$ in the argument $\omega$. Computing the real part, i.e., the PV integral is slightly more involved: By the linear interpolation performed for the imaginary part, $\delta_{b}(\omega,\epsilon)$ can be viewed as a piecewise linear function. We split the PV integral over $\delta_{b}(\omega^{\prime},\epsilon)$ into PV integrals over linear functions $(a_{i,\epsilon}\omega^{\prime}+b_{i,\epsilon})$ on intervals $\omega_{i}^{\prime},\omega_{i+1}^{\prime}]$. These are evaluated using the formula
@@
The sum over all PV integrals of the form (37) then yields the real part of the broadened Keldysh kernel $K_{b}(\omega,\epsilon)$. For this scheme to be accurate, the extent $e_{\max},e_{\max}]$ of the logarithmic $\omega^{\prime}$ grid should be significantly larger than the frequency box delimited by $\omega_{\max}$. Comparing the values of $e_{\max}$ given in Tab. 3 with our default frequency box size $\omega_{\max}=0.65$, one verifies that this is the case.
@@
[…]
@@
Figure 11: Top row (a,b): TCI ranks of $\Gamma^{\uparrow\uparrow}$ and $\Gamma_{\mathrm{core}}^{\uparrow\uparrow}$ in the three channels $a,p,t$. Bottom row (c,d): TCI ranks of $\Gamma$ and $\Gamma_{\mathrm{core}}$ in the $p$-channel for the two flavors $\uparrow\uparrow$ and $\uparrow\downarrow$. Matsubara vertices (a,c) were compressed with $\tau=10^{-3}$ and $R=10$, Keldysh vertices (b,d) with $\tau=10^{-2}$ and $R=12$.
@@ Appendix B Compression of 1D and 2D vertex contributions
[…]
@@ Appendix C Compression for different channels, flavors, Keldysh components
In this section we investigate the compressibility of core and full vertices for different flavors ($\uparrow\uparrow,\uparrow\downarrow$), frequency channels ($r=a,p,t$) and Keldysh components $\boldsymbol{k}=(k_{1}k_{2}k_{3}k_{4})$.
@@ Appendix D Frequency conventions
Appendix E Evaluating the Keldysh vertex on nonlinear grids
As mentioned in Sec. II.1, the PSFs consist of spectral peaks residing on a logarithmic energy grid. This raises the question whether the vertices could also be computed on logarithmic (or arbitrary nonlinear) grids. Since the Matsubara vertex is inherently defined on an equidistant grid, we consider this question to be relevant only for the Keldysh vertex. While the spectral representation and symmetric estimators yield the Keldysh vertex at arbitrary frequency points, evaluating the vertex on a nonlinear grid turns out to be much more expensive. This section explains why that is the case.
[…]
(i) For each individual point $\boldsymbol{\omega}\in{W}^{3}$, compute the broadened kernels upon evaluating $\Gamma^{\boldsymbol{k}}(\boldsymbol{\omega})$. This is not implemented in our code [61], but we expect it to be prohibitively expensive even if caching of previously computed kernel values were introduced.
(ii) Precompute the kernels on a very fine, equidistant grid and evaluate the vertex by interpolation from linear grids. This approach is implemented in our code [61] and has been employed in Ref. 16 when investigating the fulfillment of diagrammatic identities by mpNRG data.
In conclusion, we do not see an affordable method of evaluating the Keldysh vertex ’directly’ on a large, nonlinear grid. However, interpolation from a linear grid provides a viable alternative: Using QTCI, one can determine at which grid size the vertex is resolved to the desired accuracy and then use the resulting quantics tensor train for interpolation.
Finally, we report a preliminary investigation of the QTCI-compressibility of Keldysh vertex components stored on logarithmic grids: We compressed different Keldysh components of the $t$-channel Keldysh core vertex at $\beta=2000,\,u=0.50$ computed on a $243\times 243\times 243$ logarithmic grid. The TCI tolerance was $\tau=10^{-2}$. The grid spans a large frequency range from 3.18-3.18 to $3.18$. The vertex was interpolated trilinearly from a quantics grid with $R=15$ bits, see Ref. 16. The results are summarized in Tab. 6, showing large QTT ranks $\chi_{\mathrm{log}}$ compared to the QTT ranks on a linear grid, $\chi_{\mathrm{lin}}$. This indicates that a QTCI compression of Keldysh vertex data on a logarithmic grid is not effective.
$\boldsymbol{k}$ $1111$ $2111$ $2121$ $2112$ $1222$
$\chi_{\mathrm{log}}$ 258 475 718 379 445
$\chi_{\mathrm{lin}}$ 129 150 177 127 101
Table 6: QTT ranks of different components of the $t$-channel Keldysh core vertex $\Gamma_{\mathrm{core}}^{\uparrow\uparrow}$ computed on a logarithmic grid, and, for comparison, on an $R=15$ linear quantics grid. The tolerance is $\tau=10^{-2}$.